The President of Musicland has asked for help to evaluate the company’s performance during the current year. In particular, we are requested to calculate Musicland’s gross profit and its return on sales ratio and then explain what these measures indicate, based on the following financial information:

Net Sales $100,000

Cost of Goods Sold 45,000

Net Income 10,500
Answer #1

rate positively ..

i Net sales 100000
ii Cost of goods sold 45000
iii=i-ii Gross profit margin 55000
iv Net income 10500
v=iii/i Gross profit % = 55.00%
vi=iv/i Return on sales/Net income margin = 10.50%
Explanation:
Gross profit % tells us the gross profit which company is making on every $ of sales made
55% gross profit margin means that at every 100 USD sale company is making gross profit margin of $55.
Return on sale percent of 10.5 means that for every 100$ of sale firm is making net profit of $10.5